what equation is parallel to y=1/2x-5 that goes through the points (6,-8)
Please Help.

Found 2 solutions by mananth, scott8148:
Answer by mananth(16946)

y = 0.5x-5

Compare this equation with y=mx+b
slope m = 1/ 2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 1/ 2
m= 1/ 2 ,point (6,-8)
Find b by plugging the values of m & the point in
y=mx+b
-8 = 3 + b
b= -11
m= 1/2
Plug value of the slope and b in y = mx +b
The required equation isy = 1/2 x -11

m.ananth@hotmail.ca

Answer by scott8148(6628)

parallel means that the slope (1/2) is the same ___ what changes is the y-intercept (-5)

to find the new intercept, just plug in the new given point

y = 1/2 x + b ___ (-8) = 1/2 (6) + b ___ -11 = b

the new equation is ___ y = 1/2 x - 11